find the sum of the first 10 terms of the series 9+18+36+...
Answer by dolly(163)

Hi,
The given series 9 + 18 + 36 +.... is a geometric series whose first term(a) is 9 and the common ratio(r) is 2.
So sum upto 10 terms = a(1-r^n)/(1-r) where n = 10 here
= 9(1-2^10)/(1-2)
= 9(2^10 - 1)/1 [ multiplying by -1 on both numerator and denominator]
= 9(2^10 -1)
= 9(1024-1)
= 9(1023)
= 9207
sum upto 10 terms = 9207
Good luck!!!
dolly